OpenStudy (anonymous):

Suppose that f and g are functions such that g(-2) = 2, f(2) = 14, g'(-2) = 1, and f '(2) = -1/16. If h(x) = f(g(x)) then the equation of the line tangent to y = h(x) at x=-2 is y =

OpenStudy (dumbcow):

By the chain rule: h'(x) = f'(g(x))g'(x) h'(-2) = (-1/16)(1) h'(-2)= -1/16 slope of tangent line is -1/16 h(-2) = f(g(-2)) = 14 y-14 = -1/16(x+2) y = -1/16x +111/8